5. You own a call option on the stock with current price So = 4, the up factor u = 2. the ..down factor, d-1/2, the risk-free interest rate r-1/4 and the strike price K-5. You paid the risk-neutral price of $1.20 for this option, and you want to hedge your position (i.e. reduce your risk) so that you end up with $1.50, (as if you had invested $1.2 at the risk free rate) regardless of whether H or T occurs. Assume that you cannot sell the option, and you have no cash. You may only buy (or short sell) shares of stock or bond (borrow or invest at the risk-free interest rate), construct a portfolio to ensure you have $1.50 at t-1
Answer #1

Here given problem we had.

x0 = $ 1.2

x1 = $ 1.5

S0 = 4

u = 2

d = 1/2

r = 1/4

k = 5

t = 1

then S1(H)= USo ; S1(H)=8

So(T) = d So; S1(T) =1/2 *4,S1(T)=2

Than the hedge the option we may show with the capital x0 = $ 1.2 ,Δ0 = 1/2 shares at the time 0,and the time 0, and that investing (traingle zero s0 is 1.2 - 1 /2*2

= 1.2 - 2 = 0.8

--- At that time 1 the cash position wwould be (1+r) (Xo - Δ0) = -1,and that the stock position would be either

= 1/2 S0 (H) =4 or 1/2 S0 (T) =1

= X1(H) = 1/2 S1(H) + (1+R) X0 - Δ0 S0) =3

= 1//2 * 8 + (1 + 1/4) (0.8)

(OR)

= X1(T) =1/2 S1(T)+(1+R) X0 - Δ0 S0) =0

= 1/2*2 + (1+1/4) (-0.8)

On the other hand pay off of call was ( S1(H)--K)^T = 8-5 =3 IF W =H ;OR ( S1(T)--K)^T = (2-5)^T=0 IF W =T

As we see that the stock and that bond portfolio worth x0 = 1.2 hedges (replicate) the call . By the no arbitrage principle time - 0 value of the call option must be X0 = 1.2